Today I noticed that there was a lot of noise at the Bounce and Roll eigenfrequencies for ETMY. I found that the bandstop filter were set at completely the wrong frequencies, so I've remade them.
The filters were last tuned by Leo in May of 2011. Even so, he left the frequencies at the frequencies of the old MOS suspensions which had f_bounce ~ 12 Hz.
The FOTON plot shows the OLD ones versus the NEW ones. The DTT spectra shows the oplev error signals in the usual state. I have also copied these over to the SUSPOS,PIT,YAW, and SIDE filter banks and turned them all ON.
I also turned OFF and deleted the 3 Hz RG filter that was there. There's no such peak in the error signal and even if one wanted to compensate for the stack mode, it should be a low Q filter, not this monster. |